# Class-Incremental Learning using Diffusion Model for Distillation and Replay

### v1

@@ 4.3 Ablation Study
| āw/ SD distillation | 64.62 | 63.12 | 62.15 |
| āw/ SD distillation w/o new | 62.66 | 61.93 | 61.48 |
| āw/ SD Replay | 64.95 | 62.44 | 59.59 |
| āw/ SDDR | 65.77 | 63.84 | 61.73 |
@@
| āw/ SDDR | 64.84 | 62.71 | 60.64 |
| āw/ SDDR | 65.77 | 63.84 | 61.73 |
| āw/ SDDR | 66.41 | 64.91 | 62.73 |
| āw/ SDDR | 66.22 | 65.03 | 62.79 |
@@
Components: By leveraging synthetic images supposedly belonging to the same classes as the training dataset, it is possible to use them for both replay and distillation. TableĀ 3 compares the performance when using the synthetic images only for distillation (w/ SD distillation), only for replay (w/ SD Replay), and for both combined (w/ SDDR). Distillation and Replay using synthetic data appear to be complementary, Replay achieving higher performances for short number of incremental steps and Distillation for higher number of incremental steps. Combining both in SDDR achieves an overall improvement over both.
@@
Size of the synthetic dataset: TableĀ 4 shows the impact of the number of synthetic images generated for each past class used during training. With only 50 synthetic images per class, our method can improve the average incremental accuracy by up to 3.57p.p. compared to the standard LUCIR. This can be further increased to 5.72p.p. by expanding the size of the synthetic dataset. However, it appears that beyond a certain limit, further increasing the size of the synthetic dataset does not yield any significant improvement. This supports our approach of not using Stable Diffusion in an online manner to generate each training batch if the storage is not a constraint of the system.
@@
Size of the replay memory: TableĀ 2 measures the improvement from our method depending on the size of the replay memory containing exemplar of past classes. While the improvement is more significant in the case of a small memory size, our method also significantly improves the average incremental accuracy of both LUCIR and iCaRL when combined with a larger replay memory. Compared to PlaceboCILĀ [17], our method achieves similar performance.
@@
Use of new data: In their recent work, LiuĀ et al.Ā [17] emphasized that using data from new classes for the distillation loss was detrimental and proposed to only use data of past classes from the memory in addition to the unlabeled samples from the external additional dataset. Following their study, we also report in TableĀ 3 the results of our method while using the distillation loss only on the synthetic images of past classes and on the exemplar from the memory (w/ SD distillation w/o new). However, it appears that in our case, this significantly decreases the performance of the model. As further discussed below, we suppose that it may be due to the gap between the distribution of synthetic images and the distribution of real images: while real samples from new images are generally harmful to the distillation loss, they may be useful to transfer general knowledge about real data in our case.
@@
From synthetic to real images: As highlighted by SariyildizĀ et al.Ā [25], while the synthetic images generated by Stable Diffusion tend to be highly similar to real ones, there remains a gap that limits the generalization of the model and may negatively impact the performance. We first observed this phenomenon when using the synthetic dataset generated by Stable Diffusion as a direct replacement for the replay memory in the exemplar-free class-incremental learning setting. As depicted in TableĀ 5, when using synthetic images as exemplars for the past classes, LUCIR has difficulty generalizing to the real test dataset and achieves significantly lower accuracy compared to using real exemplars in the memory. Moreover, while it performs better than exemplar-based methods used without any replay memory, it is not competitive with approaches specifically designed for this challenging setting. Secondly, the impact of the gap between synthetic and real data may also be observed at the classifier level when using our proposed SDDR method in the most challenging settings where the synthetic images highly outnumber real exemplars in the small replay memory. Finetuning the classifier on a small dataset of real images or using an NME classifier of real exemplars only may help in this situation: for example, on CIFAR100 with 25 incremental steps and 5 exemplars per class in the replay memory, using an NME classifier for LUCIR w/ SDDR would improve the average incremental accuracy by 0.89p.p., reaching 57.21% while for standard LUCIR it would decrease it by 0.01p.p.. SariyildizĀ et al.Ā [25] addressed the sim-to-real gap by using strong augmentation policy and multi-crops and experimentally found that it results in a large performance improvement: 51% relative increase of the accuracy compared to standard limited augmentation in their simplest variant of synthetic ImageNet-100. The same approach could also be used for our method, especially as strong augmentations and larger models have been shown to be beneficial for general incremental learningĀ [14].
@@ 5 Conclusion
In this work, we showed that pretrained generative text-to-image model can be seamlessly combined with general methods for class-incremental learning in order to further improve their performances. Our approach leverages Stable Diffusion to generate labeled synthetic images belonging to the same classes as the ones previously encountered by the model and use those synthetic data for both the distillation loss and the classification loss. Complete experiments show that our approaches significantly increase the average incremental accuracy of state-of-the-art methods for class-incremental learning, especially in settings with highly restricted memory, and achieve comparable or superior results compared to other competitive approaches relying on additional real data samples.

### v2

@@ 4.3 Ablation Study
| w/ SD Distillation | 64.62 | 63.12 | 62.15 |
| w/ SD Distillation w/o new | 62.66 | 61.93 | 61.48 |
| w/ SD Replay | 64.95 | 62.44 | 59.59 |
| w/ SDDR | 65.77 | 63.84 | 61.73 |
@@
| w/ SDDR | 64.84 | 62.71 | 60.64 |
| w/ SDDR | 65.77 | 63.84 | 61.73 |
| w/ SDDR | 66.41 | 64.91 | 62.73 |
| w/ SDDR | 66.22 | 65.03 | 62.79 |
@@
Components: By leveraging synthetic images supposedly belonging to the same classes as the training dataset, it is possible to use them for both replay and distillation. Table 4 compares the performance when using the synthetic images only for distillation (w/ SD Distillation), only for replay (w/ SD Replay), and for both combined (w/ SDDR). Distillation and Replay using synthetic data appear to be complementary, Replay achieving higher performances for a short number of incremental steps and Distillation for a higher number of incremental steps. Combining both in SDDR achieves an overall improvement over each used individually.
@@
Size of the synthetic dataset: Table 5 shows the impact of the number of synthetic images generated for each past class used during training. With only 50 synthetic images per class, our method can improve the average incremental accuracy by up to 3.57p.p. compared to the standard LUCIR. This can be further increased to 5.72p.p. by expanding the size of the synthetic dataset. However, it appears that beyond a certain limit, further increasing the size of the synthetic dataset does not yield any significant improvement. This supports our approach of not using Stable Diffusion in an online manner to generate each training batch if the storage is not a constraint of the system.
@@
Size of the replay memory: Table 3 measures the improvement from our method depending on the number of incremental steps and the size of the replay memory containing exemplars of past classes. While the improvement compared to the baseline method is more significant in the case of a small memory size, our method also significantly improves the average incremental accuracy of both LUCIR and iCaRL when combined with a larger replay memory. Likewise, the improvement is more notable for more difficult settings containing numerous incremental steps. For example, when trained for 25 incremental steps with 5 samples per class in the replay memory, SDDR almost doubles the final overall accuracy for iCaRL, increasing it from 20.32% to 40.39% and augments the average incremental accuracy by 18.56p.p.. Those results highlight that our method can boost the performance of the baseline method it is combined with independently of the setting. Additional results for CIFAR100 with 5 incremental steps can be found in the Supplementary materials.
@@
Use of new data: In their recent work, Liu et al. [19] emphasized that using data from new classes for the distillation loss was detrimental and proposed to only use data of past classes from the memory in addition to the unlabeled samples from the additional external dataset. Following their study, we also report in Table 4 the results of our method while using the distillation loss only on the synthetic images of past classes and on the exemplar from the memory (w/ SD Distillation w/o new). However, it appears that in our case, this significantly decreases the performance of the model. As further discussed below, we suppose that it may be due to the gap between the distribution of synthetic images and the distribution of real images: while real samples from new images are generally harmful to the distillation loss, they may be useful for transferring general knowledge about real data in our case.
@@
From synthetic to real images: As highlighted by Sariyildiz et al. [27], while the synthetic images generated by Stable Diffusion tend to be highly similar to real ones, there remains a gap that limits the generalization of the model and may negatively impact the performance. We first observed this phenomenon when using the synthetic dataset generated by Stable Diffusion as a direct replacement for the replay memory in the exemplar-free class-incremental learning setting. As depicted in Table 6, when using synthetic images as exemplars for the past classes, LUCIR has difficulty generalizing to the real test dataset and achieves significantly lower accuracy compared to using real exemplars in the memory. Moreover, while it performs better than exemplar-based methods used without any replay memory, it is not competitive with approaches specifically designed for this challenging setting. Secondly, the impact of the gap between synthetic and real data may also be observed at the classifier level, especially when using our proposed SDDR method in the most challenging settings where the size of the replay memory is highly limited. Fixing old class embeddings learned using real data only, using an NME classifier of real exemplars only, or finetuning the classifier on a small dataset of real images may help in this situation: for example, on CIFAR100 with 25 incremental steps and 5 exemplars per class in the replay memory, using an NME classifier for LUCIR w/ SDDR would improve the average incremental accuracy by 0.89p.p., reaching 57.21% while for standard LUCIR it would decrease it by 0.01p.p.. Sariyildiz et al. [27] addressed the sim-to-real gap by using strong augmentation policy, and multi-crops and experimentally found that it results in a large performance improvement: 51% relative increase of the accuracy compared to standard limited augmentation in their simplest variant of synthetic ImageNet-100. The same approach could be used for our method, especially as strong augmentations and larger models have been shown to be beneficial for incremental learning [16].
@@ 5 Conclusion
In this work, we showed that pretrained generative text-to-image model can be seamlessly combined with general methods for class-incremental learning in order to further improve their performances. Our approach leverages Stable Diffusion to generate labeled synthetic images belonging to the same classes as the ones previously encountered by the model and use those synthetic data for both the distillation loss and the classification loss. Complete experiments show that our approaches significantly increase the average incremental accuracy of state-of-the-art methods for class-incremental learning, especially in settings with highly restricted memory, and achieve comparable or superior results compared to other competitive approaches relying on additional real data samples while offering more flexibility.
@@ Acknowledgments
Supplementary materials
1 Alternative version of ImageNet-Subset
Additionally, we considered an alternative set of classes for ImageNet-Subset, denoted as “ImageNet-Subset Alt.” in this work, used by the authors of FOSTER [32]. In Table S1, we report the performance of FOSTER combined with our method on this variation of ImageNet-Subset.
Compared to ImageNet-Subset, our proposed method SDDR achieves both higher average incremental accuracy and higher improvement over the baseline method on ImageNet-Subset Alt. .
| Methods | ImageNet-Subset Alt. | ||
| =5 | 10 | 25 | |
| FOSTER* | 78.52 | 76.49 | 71.24 |
[…]
2 Dual branch FOSTER
In our main experiments, to fairly compare our proposed method with other methods, we only evaluated the accuracy of FOSTER after performing the feature compression. To better compare with expansion-based method such as DER [35] whose number of parameters is growing at each incremental steps, we report in Table S2 the accuracy of the dual branch FOSTER model before the compression, denoted FOSTER-B4.
Experiments show that, if the memory and computational budget of the system allow it, using the dual branch FOSTER for inference with our method further improve the accuracy. On CIFAR100, FOSTER combined with SDDR achieves higher performances than DER for both single and dual branch evaluation. Furthermore, on ImageNet-Subset with 10 incremental steps, FOSTER-B4 combined with SDDR achieves a Top-1 Average incremental accuracy 0.49p.p. lower and a Top-5 Average incremental accuracy 0.50p.p. higher than DER while requiring about five times less parameters.
| Methods | CIFAR100 | ImageNet-Subset | ||||
| =5 | 10 | 25 | 5 | 10 | 25 | |
| DER (w/o Pruning) [35] | 68.52 | 67.09 | - | - | 78.20 | - |
| FOSTER* [32] | 71.17 | 68.89 | 65.07 | 76.09 | 75.05 | 70.96 |
[…]
| FOSTER-B4* [32] | 72.08 | 69.32 | 65.41 | 77.48 | 76.11 | 71.40 |
[…]
3 CIFAR-100 with 5 incremental steps
Following Liu et al. [19], we report in Table S3 additional results for the different baseline methods on CIFAR100 with a base step containing half of the classes followed by 5 incremental steps depending on the number of exemplars saved in memory for each class.
When FOSTER [32] is used with a really small replay memory, we observed that the improvement resulting from using our proposed method SDDR is limited. We suppose that this is due to the logits alignment loss used by the authors that too strongly limits the plasticity of the model and restricts our proposed method. By tuning the hyperparameters of the logits aligment loss, it should be possible to further improve the performance of FOSTER when combined with our proposed method.
| Methods | CIFAR100 with =5 | |||||||
| 5 exemplars/class | 10 exemplars/class | 20 exemplars/class | 50 exemplars/class | |||||
| Average | Last | Average | Last | Average | Last | Average | Last | |
| iCaRL* [24] | 43.40 | 31.90 | 52.30 | 40.11 | 57.68 | 47.45 | 62.09 | 54.23 |
| w/ PlaceboCIL [19] | 51.55 | 39.35 | 59.11 | 46.42 | 61.24 | 51.47 | - | - |
[…]
| LUCIR* [10] | 53.14 | 41.52 | 60.77 | 49.65 | 63.37 | 53.91 | 65.63 | 57.58 |
| w/ PlaceboCIL [19] | 62.74 | 53.25 | 64.79 | 55.44 | 65.28 | 56.23 | - | - |
[…]
| FOSTER* [32] | 56.94 | 43.74 | 63.09 | 54.34 | 71.17 | 63.97 | 70.00 | 64.10 |
| w/ PlaceboCIL [19] | 62.78 | 50.72 | 65.12 | 54.81 | 71.97 | 64.43 | - | - |
[…]